A rocket is 80m long on the ground.when it is in flight,it's length is 99mm to an observer on the ground. Calculate it's speed(1.77×10^8m/s)


1
Expert's answer
2021-07-13T10:42:44-0400
"l=l_0\sqrt{1-\frac{v^2}{c^2}}\\0.099=80\sqrt{1-\frac{v^2}{(3\cdot10^8)^2}}\\v=2.99792228\cdot10^8\frac{m}{s}"